OpenStudy (anonymous):

what is the parallel to 2x-3y=4 passing through (2,6)

OpenStudy (marigirl):

first i am going to rewrite the equation in the form y=mx+c 2x-3y=4 -3y=4-2x y=-4/3+2/3x \[y=\frac{ 2 }{ 3 }x-\frac{ 4 }{ 3 }\]

OpenStudy (marigirl):

from this we can see that the gradient of that equation is m=2/3

OpenStudy (marigirl):

If this new line is parallel to this equation then it will have the same gradient.. therefore m=2/3 .. and from the question we have a given coordinate (2,6)

OpenStudy (marigirl):

what we can do now is use this information to obtain that new equation y-y1=m(x-x1) y-6 =(2/3)(x-2) expand \[y-6=\frac{ 2 }{ 3 }x-\frac{ 4 }{ 3 }\] then add 6 t bth sides \[y=\frac{ 2 }{ 3 }x-\frac{ 22 }{ 3 }\]

OpenStudy (anonymous):

Marigirl, many thanks for your help.

OpenStudy (marigirl):

no prbs! can u plz give me a medal